### v1

@@ Abstract
Modeling and calibrating the fidelity of synthetic data is paramount in shaping the future of safe and reliable self-driving technology by offering a cost-effective and scalable alternative to real-world data collection. We focus on its role in safety-critical applications, introducing four types of instance-level fidelity that go beyond mere visual input characteristics. The aim is to align synthetic data with real-world safety issues. We suggest an optimization method to refine the synthetic data generator, reducing fidelity gaps identified by the DNN-based component. Our findings show this tuning enhances the correlation between safety-critical errors in synthetic and real images.
Keywords:
AI safety fidelity optimization synthetic data
1 Introduction
@@
The increasing reliance on synthetic data for virtual testing is crucial in solving contemporary engineering challenges. In the automotive sector, scenario-based virtual testing with synthetic data is now the norm for ensuring autonomous vehicle (AV) [10, 1] safety, replicating real-world scenarios. Significant investments are made in developing advanced simulators [3, 16] and Generative AI (GAI) [24, 8] models to improve aspects like fidelity, creating realistic virtual environments and sensors that closely resemble real driving conditions. However, achieving perfect fidelity in synthetic data generation remains a challenge. This leads to a critical question driving our research: “What level of fidelity is necessary for synthetic data to be deemed adequate for safety purposes?
@@
In this paper, we focus on “instance-level fidelity”, a term we introduce to describe how accurately an individual synthetic data point, such as an image, replicates the real world’s characteristics, features, and variability. We provide a mathematical framework for four different fidelity types, illustrated in Fig. 1. At the left end of this spectrum is “Input Value fidelity” (IV-fidelity), which requires pixel-level similarity between synthetic and real-world images. IV-fidelity is model-agnostic, aiming for a universally high fidelity level suitable for any System-Under-Testing (SUT). Despite the stringent criteria of IV-fidelity, it does not guarantee the sameness of output, mainly due to Deep Neural Networks’ (DNN) lack of robustness [5]. Therefore, it’s more effective to evaluate fidelity by considering how the specific SUT processes synthetic data and its intended use, rather than solely focusing on strict fidelity requirements.
@@
Towards this direction, we thus introduce the fidelity from the restricted view using the safety perspective, i.e., the “Safety-Aware fidelity” (SA-fidelity) on the rightmost of Fig. 1. Conceptually, what differentiates a superior synthetic data generator of higher SA-fidelity from a normal one is: the capability to generate safety-critical inputs that can, when interpreting the semantics of the input and reconstructing its scenario in the real world, lead to similar safety-critical concerns. For example, consider a scenario where a pedestrian is extremely far from the vehicle. Even when the synthetic image generator fails to produce the appropriate contour for the distant pedestrian in a 2D image plane due to the object being extremely small, there exist no safety concerns, as whether or not a DNN-based perception system detects the pedestrian at the moment does not matter. This ultimately leads to a pragmatic, risk-aware definition of instance-level fidelity that does not demand unnecessary perfection, such as pixel-level precision between synthetic and real images.
@@
With fidelity being a SUT-specific measure, we further consider techniques to adjust the synthetic data generation process to increase the SA-fidelity. Classical synthetic data generators may have parameters such as noise or camera angle, while neural synthetic data generators have learnable parameters. These parameters can be further calibrated accordingly to fit the specific need. The calibration of the synthetic data generator can thus be viewed as an optimization problem to reduce the discrepancy in SA-fidelity, where techniques such as genetic programming or gradient descent can be applied depending on the type of synthetic data generation.
@@
In summary, the main contribution of this paper includes the following:
@@
2 Foundations
@@
Denote $\mathbb{A}$ as the set of semantic attributes (e.g., moving and scenery objects) that can be characterized by a scenario description language. Without loss of generality (WLOG), we assume each semantic attribute $a\in\mathbb{A}$ is a parameter that takes on a real number [^1] Other types of attributes, e.g., categorical ones, can be encoded as real numbers. as its value. A scenario description $sd=\{(a_{1},v_{1}),\dots,(a_{k},v_{k})\}$ contains a subset of $k$ attributes $\{a_{1},\ldots,a_{k}\}\subset\mathbb{A}$ and associates each attribute with a concrete value $v_{i},i\in[1..k]$.
@@
Let $x\in\mathbb{R}^{d_{0}}$ be a data point (image, lidar point-cloud) with $d_{0}$ being its dimension. A synthetic data generator $g_{\theta}$ takes a scenario description $sd$ and creates a data point $x^{s}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}g_{\theta}(sd)\in\mathbb{R}^{d_{0}}$; $g_{\theta}(sd)$ can either be deterministic or stochastic, whereby calling $g_{\theta}$ stochastic, we refer to the situation that $g_{\theta}(sd)$ may not generate the same results due to sampling. The parameter $\theta$ denotes configurable parameters of the synthetic data generator that normally will not be described by scenario description languages, e.g., the virtual sensor settings in simulators and the guidance scale setting in GAI models. Given a scenario description $sd$, let ${\textsf{{RW}}}(sd)\subseteq\mathbb{R}^{d_{0}}$ be an oracle function that returns the infinite set of real-world data points whose semantic attributes match that of $sd$. Note that given a scenario description $sd$, the images produced by the synthetic data generator may not be strictly contained in ${\textsf{{RW}}}(sd)$, resulting in fidelity concerns.
@@
Let $f:\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}^{d_{L}}$ be a perception function that takes an input data point $x\in\mathbb{R}^{d_{0}}$ and produces an output vector of $d_{L}$ dimensions. Consequently, $f$ can accept an input data point $x^{s}$ taken from $g_{\theta}(sd)$ as well as any data point within ${\textsf{{RW}}}(sd)$. When $f$ is a DNN, we use $f^{l}(x)$ to represent the feature vector of $f$ at the $l$-th layer under the input being $x$.
3 The Facets of Instance-level Fidelity
@@
3.1 Input, output, and latent feature fidelity
@@
A synthetic data point $x^{s}\in g_{\theta}(sd)$ is defined to be $\langle\mathcal{D}_{in},\epsilon_{in}\rangle$ Input-Value (IV) fidelitous iff
$\exists\tilde{x}\in\emph{{{{RW}}}}(sd):\mathcal{D}_{in}(x^{s},\tilde{x})\leq\epsilon_{in}$ (1)
@@
$\exists\tilde{x}\in\emph{{{{RW}}}}(sd):\mathcal{D}_{out}(f(x^{s}),f(\tilde{x}))\leq\epsilon_{out}$ (2)
@@
The following result connects IV-fidelity, OV-fidelity, and the standard definition of local robustness of DNNs.
@@
Lemma 1 enables us to build an instance-level interpretation regarding the slang “photo-realistic is not real” heard in the field of autonomous driving, and its raised safety concerns: The phrase “photo-realistic” can be interpreted as a synthetic input $x^{s}$ being IV-fidelitous, while the phrase “is not real” reflects the concern that under the processing of the DNN, the real and the synthetic image may not lead to the sameness of output, i.e., OV-fidelity can not be guaranteed. This is analogous to adversarial examples [5], where OV-fidelity does not hold despite IV-fidelity may hold, reflecting the differences between the perturbed and the original image being human-eye indistinguishable. Altogether, one shall rephrase the slang to “photo-realistic is guaranteed to be real only when the difference between pixels is within the local robustness bound”.
Finally, apart from fidelity manifested in the input or output values, one can also define fidelity based on the sameness of the learned features between synthetic and real-world images, as demonstrated below. The sameness of learned features hints at the sameness of the decision mechanism, but again, analogous to the proof of Lemma 1, one can also formally deduce that LF-fidelity does not immediately imply OV-fidelity.
@@
$\exists\tilde{x}\in\emph{{{{RW}}}}(sd):\;\forall l\in\mathcal{L}_{lf}:\mathcal{D}_{lf}(f^{l}(x^{s}),f^{l}(\tilde{x}))\leq\epsilon_{lf}$ (3)
@@
3.2 Safety-aware fidelity
Before delving into technical definitions, we first explain the design intuition of SA-fidelity. When applying DNN $f$ with two data points $x^{s}$ and $x^{\prime}$, the meaning of (output prediction) SA-fidelity $f(x)\equiv_{s}f(x^{\prime})$ enables characterizing the behavior that both DNN output predictions do not lead to inconsistent safety concerns.
@@ 3.2 Safety-aware fidelity
Recall that scenario description $sd=\{(a_{1},v_{1}),\dots,(a_{k},v_{k})\}$ contains a subset of $k$ attributes $a_{1},\ldots,a_{k}\in\mathbb{A}$ and associates each attribute with a concrete value. Let prediction influencing attributes ${\textsf{{pia}}}(sd)$ be a subset of a scenario description $sd$, where the attributes may be perceived by the perception unit. Let safety influencing attributes ${\textsf{{sia}}}(sd)$ return a subset of prediction influencing attributes ${\textsf{{pia}}}(sd)$, where an incorrect perception of these attributes can lead to safety concerns.
@@
Consider a perception unit performing 3D object detection on cars for automatic emergency braking (AEB), and consider a scenario where there are front cars, one close and one distant to the autonomous vehicle, in rainy situations. The scenario description can be $sd\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{({\textsf{{frontcar}}},1),({\textsf{{farcar}}},1),\dots,({\textsf{{rain}}},1)\}$, where the semantic attribute “frontcar” indicates the existence of front vehicles being close-by (subject to a predefined physical distance), “farcar” indicates the existence of distant vehicles, and “rain” indicates if the scenario is rainy. The ${\textsf{{pia}}}(sd)$ and ${\textsf{{sia}}}(sd)$ may act as follows.
  • ${\textsf{{pia}}}(sd)=\{({\textsf{{frontcar}}},1),({\textsf{{farcar}}},1)\}$, as the perception unit only infers knowledge on the existence of vehicles.
  • ${\textsf{{sia}}}(sd)=\{({\textsf{{frontcar}}},1)\}$, due to distant cars having no safety concerns.
Subsequently, let $x\in\{g_{\theta}(sd)\}\cup{\textsf{{RW}}}(sd)$, and define ${\textsf{{inter}}}(x,f,sd)$ be an interpretation function that “rewrites” the DNN output vector $f(x)$ into a scenario description format, by changing ${\textsf{{pia}}}(sd)$ and keep the rest the same as that of $sd$, due to DNN unable to recognize semantic attributes within $sd\setminus{\textsf{{pia}}}(sd)$.
@@
Consider again the scenario description $sd\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{({\textsf{{frontcar}}},1),\\ ({\textsf{{farcar}}},1),\dots,({\textsf{{rain}}},1)\}$, where for an input $x\in{\textsf{{RW}}}(sd)$, if the prediction $f(x)$ fails to identify any vehicle, then
${\textsf{{inter}}}(x,f,sd)=\{({\textsf{{frontcar}}},0),({\textsf{{farcar}}},0),\dots,({\textsf{{rain}}},1)\}$
@@
Given scenario description $sd$, function $f$, and two input data points $x$ and $x^{\prime}$ where $x,x^{\prime}\in\{g_{\theta}(sd)\}\cup\emph{{{{RW}}}}(sd)$. The output safety similarity $f(x)\equiv_{s}f(x^{\prime})$ holds iff the following condition holds.
[…]
@@
For an input $x^{\prime}\in{\textsf{{RW}}}(sd)$, if $f(x^{\prime})$ fails to identify any vehicle, but for the synthetic data point $x^{s}=g_{\theta}(sd)$, $f(x^{s})$ only fails to identify any far-distant vehicle. Then $f(x^{s})\not\equiv_{s}f(x^{\prime})$ due to the following reasoning:
${\textsf{{inter}}}(x^{\prime},f,sd)=\{({\textsf{{frontcar}}},0),({\textsf{{farcar}}},0),\dots,({\textsf{{rain}}},1)\}$
${\textsf{{inter}}}(x^{s},f,sd)=\{({\textsf{{frontcar}}},1),({\textsf{{farcar}}},0),\dots,({\textsf{{rain}}},1)\}$
${\textsf{{sia}}}({\textsf{{inter}}}(x^{\prime},f,sd))=\{({\textsf{{frontcar}}},0)\}\neq{\textsf{{sia}}}({\textsf{{inter}}}(x^{s},f,sd))=\{({\textsf{{frontcar}}},1)\}$
@@
Given $f$, a synthetic data point $x^{s}\in g_{\theta}(sd)$ is defined to be Safety-Aware (SA) fidelitous under $\emph{{{{inter}}}}(\cdot)$ and $\emph{{{{sia}}}}(\cdot)$ iff
$\exists\tilde{x}\in\emph{{{{RW}}}}(sd):f(x^{s})\equiv_{s}f(\tilde{x})$ (5)
4 Calibrating Data Generation for Safety-aware Fidelity

### v2

@@ Abstract
Modeling and calibrating the fidelity of synthetic data is paramount in shaping the future of safe and reliable self-driving technology by offering a cost-effective and scalable alternative to real-world data collection. We focus on its role in safety-critical applications, introducing four types of instance-level fidelity that go beyond mere visual input characteristics. The aim is to ensure that applying testing on synthetic data can reveal real-world safety issues, and the absence of safety-critical issues when testing under synthetic data can provide a strong safety guarantee in real-world behavior. We suggest an optimization method to refine the synthetic data generator, reducing fidelity gaps identified by deep learning components. Experiments show this tuning enhances the correlation between safety-critical errors in synthetic and real data.
I Introduction
@@
The increasing reliance on synthetic data for virtual testing is crucial in solving contemporary engineering challenges. In the automotive sector, scenario-based virtual testing with synthetic data is now the norm for ensuring autonomous vehicle (AV) safety [1, 2], replicating real-world scenarios. Significant investments are made in developing advanced simulators [3, 4] and Generative AI (GAI) [5, 6] models to improve aspects such as fidelity, creating realistic virtual environments and sensors that closely resemble real driving conditions. However, achieving perfect fidelity in synthetic data generation remains a challenge. This leads to the question driving our work: “What level of fidelity is necessary for synthetic data to be deemed adequate for safety purposes?
@@
In this paper, we focus on “instance-level fidelity [^1] In virtual testing for AVs, fidelity may also refer to the closeness of synthetic data distribution to the real-world data distribution, and the accuracy of the sequential dynamics as the AV controller interacts with the virtual environments. While those perspectives are important, this paper focuses on examining fidelity at the level of an individual synthetic data-point, e.g., an image or a point-cloud depending on the type of sensors simulated; in this paper, we study synthetic images without loss of generality., a term we introduce to describe how accurately an individual synthetic data point, such as an image, replicates the real world’s characteristics, features, and variability. We provide a mathematical framework for four different fidelity types, illustrated in Fig. 1. At the left end of this spectrum is “Input Value fidelity” (IV-fidelity), which requires pixel-level similarity between synthetic and real-world images. IV-fidelity is model-agnostic, aiming for a universally high fidelity level suitable for any System-Under-Testing (SUT). Despite the stringent criteria of IV-fidelity, it does not guarantee the sameness of output, mainly due to Deep Neural Networks’ (DNN) lack of robustness [7]. Therefore, it’s more effective to evaluate fidelity by considering how the specific SUT processes synthetic data and its intended use, rather than solely focusing on strict fidelity requirements.
@@
Towards this direction, we thus introduce the fidelity from the restricted view using the safety perspective, i.e., the “Safety-Aware fidelity” (SA-fidelity) on the rightmost of Fig. 1. Conceptually, what differentiates a superior synthetic data generator of higher SA-fidelity from a normal one is the alignment of reproducing safety-critical events. This implies the capability to generate input with objects (e.g., pedestrians) that can, when interpreting the semantics of the input and reconstructing its scenario in the real world, lead to similar safety-critical concerns. At the same time, the absence of safety-critical issues when testing using synthetic data can provide a strong safety guarantee in real-world behavior. For example, consider a scenario where a pedestrian is extremely far from the autonomous vehicle. Even when the synthetic image generator fails to produce the appropriate contour for the distant pedestrian in a 2D image plane due to the object being extremely small, there exist no safety concerns, as whether or not a DNN-based perception system detects the pedestrian at the moment does not matter. This ultimately leads to a pragmatic, risk-aware definition of instance-level fidelity that does not demand unnecessary perfection, such as pixel-level precision between synthetic and real images.
@@
With fidelity being an SUT-specific measure, we further consider techniques to adjust the synthetic data generation process to increase the SA-fidelity. Classical synthetic data generators may have parameters such as noise or camera angle, while neural synthetic data generators have learnable parameters. These parameters can be further calibrated according to the specific need. The calibration of the synthetic data generator can thus be viewed as an optimization problem to reduce the discrepancy in SA-fidelity, where techniques such as genetic programming or gradient descent can be applied depending on the type of synthetic data generation.
@@
In summary, the main contribution of this paper includes:
@@
The paper is structured as follows. After summarizing related work in Sec. II and building the preliminaries in Sec. III, we introduce three fidelity concepts, their relations, and the extension of SA-fidelity in Sec. IV. We then present an optimization-based approach for post-calibrating the synthetic data generator and the evaluation in Sec. V and VI. Finally, we conclude with future directions in Sec. VII.
II Related Work
Simulation-based testing for autonomous systems has been an active field with accessible commercial tools. Testing non-perception modules is relatively mature (see [9] for a review), where the scenario description can be translated into objects (as bounding boxes) and their physical characteristics (e.g., velocity, coordinate), serving as the input to the motion planner. The fidelity of cyber-physical system simulation (with differential equation included) has been addressed in various frameworks such as Ptolemy II [10], Modellica [11, 12], and Mathworks Simscape [^2] https://www.mathworks.com/products/simscape.html. Our paper primarily focuses on image-based perception; simulation tools such as NVIDIA DriveSim [13] can create “photo-realistic images” as the underlying pipelines supporting physics-based rendering. These efforts lead to synthetic datasets like SynPeDS [14] tailored for pedestrian detection tasks in urban scenes. With a virtual world modeled, varying scene parameters can reveal performance-limiting factors in perception models [15]. Apart from physic-based synthetic data generation, the use of GAI also hints at strong promises [5, 6]. Despite the aforementioned progress being made, fundamental concerns remain regarding the validity of using virtual testing as a strong safety argument (i.e., to replace “real miles” with “virtual miles” [16]). This motivates us to introduce a novel perspective to existing fidelity taxonomies (e.g., [17]), and mathematically precise the meaning of fidelity by aligning it with its intended use, while SA-fidelity provides a rigorous statement to avoid over-perfecting every detail in simulation like the effort in [18].
Within the ML community, the technique of domain adaptation (see [19] for an extensive survey) allows a DNN to perform well on new input domains via techniques such as training the DNN to encourage domain confusion in the feature space. Practically, one can utilize the concept of domain adaption, leading to training with both real and synthetic data included (e.g., [20] on mixing Cityscapes (real) and GTA5 (synthetic)), thereby encouraging a DNN have consistent performance in both real and synthetic images. Our result of SA-fidelity complements these results, as 1) we enable a safety-oriented, task-specific assessment of the synthetic data (i.e., not on the DNN), and 2) our calibration on the synthetic data generator allows the DNN to remain unmodified, thereby acting on top of any existing domain adaptation scheme.
III Foundations
@@
Denote $\mathbb{A}$ as the set of semantic attributes (e.g., moving and scenery objects) that can be characterized by a scenario description language. Without loss of generality (WLOG), we assume that each semantic attribute $a\in\mathbb{A}$ is a parameter that takes on a real number [^3] Other attribute types can be encoded as real numbers. as its value. A scenario description $sd=\{(a_{1},v_{1}),\dots,(a_{k},v_{k})\}$ contains a subset of $k$ attributes $\{a_{1},\ldots,a_{k}\}\subset\mathbb{A}$ and associates each attribute with a concrete value $v_{i},i\in[1..k]$.
@@
Let $x\in\mathbb{R}^{d_{0}}$ be a data point (image, lidar point-cloud) with $d_{0}$ being its dimension. A synthetic data generator $g_{\theta}$ takes a scenario description $sd$ and creates a data point $x^{s}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}g_{\theta}(sd)\in\mathbb{R}^{d_{0}}$; $g_{\theta}(sd)$ can either be deterministic or stochastic, whereby calling $g_{\theta}$ stochastic, we refer to the situation that $g_{\theta}(sd)$ may not generate the same results due to sampling. The parameter $\theta$ denotes configurable parameters of the synthetic data generator that will normally not be described by scenario description languages, e.g., the virtual sensor settings in simulators and the guidance scale setting in GAI models. Given a scenario description $sd$, let ${{\small\textsf{{RW}}}}(sd)\subseteq\mathbb{R}^{d_{0}}$ be an oracle function that returns an infinite set of real-world data points whose semantic attributes match that of $sd$. A fidelity concern arises because the images generated by the synthetic data generator, given a scenario description $sd$, may not be strictly contained in ${{\small\textsf{{RW}}}}(sd)$.
Remark 1 (Approximation of $\emph{{{\small{{RW}}}}}(sd)$)
Given a scenario description $sd$, ${{\small\textsf{{RW}}}}(sd)$ is infinite and thus unknown in theory, which can only be approximated in practice, e.g., from a collected real-world dataset $X^{cr}\subset\mathbb{R}^{d_{0}}$. This leads to an approximated set $\widehat{{{\small\textsf{{RW}}}}}(sd)\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{x\mid x\in X^{cr}\cap{{\small\textsf{{RW}}}}(sd)\}$.
@@
Let $f:\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}^{d_{L}}$ be a perception function that takes an input data point $x\in\mathbb{R}^{d_{0}}$ and produces an output vector of $d_{L}$ dimensions. Consequently, $f$ can accept an input data point $x^{s}$ taken from $g_{\theta}(sd)$ as well as any data point within ${{\small\textsf{{RW}}}}(sd)$. When $f$ is a DNN, we use $f^{l}(x)$ to represent the feature vector of $f$ at the $l$-th layer taking input $x$.
IV The Facets of Instance-level Fidelity
@@
IV-A Input, output, and latent feature fidelity
@@
A synthetic data $x^{s}\in g_{\theta}(sd)$ is defined to be $\langle\mathcal{D}_{in},\epsilon_{in}\rangle$ Input-Value (IV) fidelitous iff
$\exists\tilde{x}\in{{\small\textsf{{RW}}}}(sd):\mathcal{D}_{in}(x^{s},\tilde{x})\leq\epsilon_{in}$ (1)
@@
$\exists\tilde{x}\in{{\small\textsf{{RW}}}}(sd):\mathcal{D}_{out}(f(x^{s}),f(\tilde{x}))\leq\epsilon_{out}$ (2)
@@
Intuitively, IV and OV-fidelity are defined based on some distance metrics between the given synthetic data and any real-world data that sharing the same $sd$, before and after being processed by the DNN $f$.
Remark 2 (Conservative bound on distance-based fidelity)
As per Remark 1, we can only implement the approximated $\widehat{{{\small\textsf{{RW}}}}}(sd)$ when measuring IV/OV-fidelity in practice. Using the IV-fidelity as an example, since apparently $\min\{\mathcal{D}_{in}(x^{s},\tilde{x})\mid\tilde{x}\in\widehat{{{\small\textsf{{RW}}}}}(sd)\}\geq\min\{\mathcal{D}_{in}(x^{s},\tilde{x})\mid\tilde{x}\in{{\small\textsf{{RW}}}}(sd)\}$, the practical approximation is calculating an upper bound on the fidelity discrepancy distance (similar reasoning holds for OV-fidelity). Thus, we may take such conservative approach that errs on the side of caution to prioritize safety.
The following result connects IV-fidelity, OV-fidelity, and the common definition of local robustness of DNNs [24].
@@
Lemma 1 enables us to build an instance-level interpretation on the jargon “photo-realistic is not real”, reflecting the common concern on the validity of using photo-realistic simulation in autonomous driving: The phrase “photo-realistic” can be interpreted as a synthetic input $x^{s}$ being IV-fidelitous, while the phrase “is not real” reflects the issue that under the processing of the DNN, the real and the synthetic image may not lead to the sameness of output, i.e., OV-fidelity can not be guaranteed. This is analogous to adversarial examples [7], where OV-fidelity does not hold despite IV-fidelity may hold, reflecting the differences between the perturbed and the original image being human-eye indistinguishable. Altogether, one shall rephrase the slang to “photo-realistic is guaranteed to be real only when the difference between pixels is within the local robustness bound”.
Finally, apart from the fidelity manifested in the input or output values, one can also define the fidelity based on the sameness of the learnt features between synthetic and real-world images, as demonstrated below. The sameness of the learnt features hints at the sameness of the decision mechanism, but again, analogously to the proof of Lemma 1, one can also formally deduce that LF-fidelity does not immediately imply OV-fidelity.
@@
$\exists\tilde{x}\in{{\small\textsf{{RW}}}}(sd):\;\forall l\in\mathcal{L}_{lf}:\mathcal{D}_{lf}(f^{l}(x^{s}),f^{l}(\tilde{x}))\leq\epsilon_{lf}$ (3)
@@
Notably, Remark 2 also holds for LF-fidelity.
IV-B Safety-aware fidelity
Prior to technical definitions, we first explain the design intuition of SA-fidelity. When applying DNN $f$ with two data points $x^{s}$ and $x^{\prime}$, the meaning of (output prediction) SA-fidelity $f(x)\equiv_{s}f(x^{\prime})$ enables characterizing the behavior that both DNN output predictions do not lead to inconsistent safety concerns.
@@ IV-B Safety-aware fidelity
Recall that scenario description $sd=\{(a_{1},v_{1}),\dots,(a_{k},v_{k})\}$ contains a subset of $k$ attributes $a_{1},\ldots,a_{k}\in\mathbb{A}$ and associates each attribute with a concrete value. Let prediction influencing attributes ${{\small\textsf{{pia}}}}(sd)$ be a subset of a scenario description $sd$, where the attributes may be perceived by the perception unit. Let safety influencing attributes ${{\small\textsf{{sia}}}}(sd)$ return a subset of prediction influencing attributes ${{\small\textsf{{pia}}}}(sd)$, where an incorrect perception of these attributes can lead to safety concerns.
@@
Consider a perception unit performing 3D object detection on cars for automatic emergency braking (AEB), and consider a scenario where there are front cars, one close and one distant to the autonomous vehicle, in rainy situations. The scenario description can be $sd\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{({{\small\textsf{{frontcar}}}},1),({{\small\textsf{{farcar}}}},1),\dots,({{\small\textsf{{rain}}}},1)\}$, where the semantic attribute “frontcar” indicates the existence of front vehicles being close-by (subject to a predefined physical distance), “farcar” indicates the existence of distant vehicles, and “rain” indicates if the scenario is rainy. The ${{\small\textsf{{pia}}}}(sd)$ and ${{\small\textsf{{sia}}}}(sd)$ may act as follows.
  • ${{\small\textsf{{pia}}}}(sd)=\{({{\small\textsf{{frontcar}}}},1),({{\small\textsf{{farcar}}}},1)\}$, as the perception unit only infers knowledge on the existence of vehicles.
  • ${{\small\textsf{{sia}}}}(sd)=\{({{\small\textsf{{frontcar}}}},1)\}$, due to distant cars having no safety concerns.
Subsequently, let $x\in\{g_{\theta}(sd)\}\cup{{\small\textsf{{RW}}}}(sd)$, and define ${{\small\textsf{{inter}}}}(x,f,sd)$ be an interpretation function that “rewrites” the DNN output vector $f(x)$ into a scenario description format, by changing ${{\small\textsf{{pia}}}}(sd)$ and keep the rest the same as that of $sd$, due to DNN unable to recognize semantic attributes within $sd\setminus{{\small\textsf{{pia}}}}(sd)$.
@@
Consider again the scenario description $sd\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{({{\small\textsf{{frontcar}}}},1),({{\small\textsf{{farcar}}}},1),\dots,({{\small\textsf{{rain}}}},1)\}$, where for an input $x\in{{\small\textsf{{RW}}}}(sd)$, if the prediction $f(x)$ fails to identify any vehicle, then
${{\small\textsf{{inter}}}}(x,f,sd)=\{({{\small\textsf{{frontcar}}}},0),({{\small\textsf{{farcar}}}},0),\dots,({{\small\textsf{{rain}}}},1)\}$
@@
Given scenario description $sd$, DNN $f$, and input data points $x$ and $x^{\prime}$ where $x,x^{\prime}\in\{g_{\theta}(sd)\}\cup{{\small\textsf{{RW}}}}(sd)$. The output safety similarity $f(x)\equiv_{s}f(x^{\prime})$ holds iff the following condition holds.
${{\small\textsf{{sia}}}}({{\small\textsf{{inter}}}}(x,f,sd))={{\small\textsf{{sia}}}}({{\small\textsf{{inter}}}}(x^{\prime},f,sd))$ (4)
@@
For an input $x^{\prime}\in{{\small\textsf{{RW}}}}(sd)$, if $f(x^{\prime})$ fails to identify any vehicle, but for the synthetic data point $x^{s}=g_{\theta}(sd)$, $f(x^{s})$ only fails to identify any far-distant vehicle. Then $f(x^{s})\not\equiv_{s}f(x^{\prime})$ due to the following reasoning:
${{\small\textsf{{inter}}}}(x^{\prime},f,sd)$ =$=\{({{\small\textsf{{frontcar}}}},0),({{\small\textsf{{farcar}}}},0),\dots,({{\small\textsf{{rain}}}},1)\}$
${{\small\textsf{{inter}}}}(x^{s},f,sd)$ =$=\{({{\small\textsf{{frontcar}}}},1),({{\small\textsf{{farcar}}}},0),\dots,({{\small\textsf{{rain}}}},1)\}$
${{\small\textsf{{sia}}}}({{\small\textsf{{inter}}}}(x^{\prime},f,sd))$ =$=\{({{\small\textsf{{frontcar}}}},0)\}$
$\neq{{\small\textsf{{sia}}}}({{\small\textsf{{inter}}}}(x^{s},f,sd))=\{({{\small\textsf{{frontcar}}}},1)\}$
@@
Given $f$, a synthetic data point $x^{s}\in g_{\theta}(sd)$ is defined to be Safety-Aware (SA) fidelitous under $\emph{{{\small{{inter}}}}}(\cdot)$ and $\emph{{{\small{{sia}}}}}(\cdot)$ iff
$\exists\tilde{x}\in{{\small\textsf{{RW}}}}(sd):f(x^{s})\equiv_{s}f(\tilde{x})$ (5)
Note, aforementioned remarks on approximating ${{\small\textsf{{RW}}}}(sd)$ and yielding conservative measurements holds for SA-fidelity.
V Calibrating Data Generation for Safety-aware Fidelity